# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

### v1

@@ Abstract
Figure 1: Disordered nonequilibrium spin model with 1000 spins. (a) Steady-state EP estimation for different inverse temperatures $\beta$. (b) Asymmetry of inferred parameters, shown against the true coupling asymmetries in the model for $\beta=2.5$ ($R^{2}=0.9831$). Estimates are based on $10^{9}$ state transitions sampled by Monte Carlo.
We consider the diluted version of the model [47, 48, 49] with $k$ average neighbors. Here $w_{ij}=c_{ij}z_{ij}/\sqrt{k}$, with binary connections $c_{ij}\sim\textrm{Bernoulli}(k/(N-1))$ and real-valued weights $z_{ij}\sim\mathcal{N}(0,1)$. Our observables of interest are time-lagged correlations,
@@
with conjugate parameters $\theta_{ij}$. Because the system has multipartite dynamics and observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes. This dramatically improves the performance scaling of our optimization in system size.
Fig. 1(a) shows the actual and inferred EP at different $\beta$ for $N=1000$, $k=6$. EP increases with $\beta$, and all estimators (${\Sigma}_{\bm{g}}^{\textrm{TUR}}$, $\widehat{{\Sigma}}_{\bm{g}}$, ${\Sigma}_{\bm{g}}$) agree in the near-equilibrium regime of small $\beta$. Moreover, ${\Sigma}_{\bm{g}}$ provides a tight bound on EP even in the far-from-equilibrium regime of large $\beta$. The gap between ${\Sigma}_{\bm{g}}$ and $\widehat{{\Sigma}}_{\bm{g}}$ indicates onset of highly non-Gaussian statistics from $\beta\approx 2$.
As discussed near Eq. (6), ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma(\vec{{\bm{x}}})$ can be closely approximated by a linear combination of observables $g_{ij}$. This is a nontrivial result even for this model, since $\sigma(\vec{{\bm{x}}})$ contains contributions from the difference of steady-state probabilities, which is not known in closed form. As we show in the SM, our optimal parameters ${\bm{\theta}}^{*}$ can be used to infer the asymmetry of the coupling constants as $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider a different family of Gaussian coupling matrices, for which this model exhibits a nonequilibrium phase transition [6]. We show that our approach accurately infers EP in this regime, which has been shown to be difficult for classical approximations [50].
@@
Figure 2: (a) Estimated EP per expected number of spikes $R$, in the Neuropixels Visual Behavior dataset for three conditions. Error bars indicate standard error of the mean. (b) Example of the values of inferred coupling parameters $\theta^{*}_{ij}$ from one trial in the active condition selecting the 200 most active neurons, grouped by visual brain area. To improve visualization, lower triangle shows $\theta_{ji}^{*}\equiv-\theta_{ij}^{*}$ for $i<j$.
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [44], containing spike-train recordings from 81 mice across 103 sessions. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures. We analyze data from visual areas during three conditions: (1) active behavior (visual change detection task), (2) passive replay (same stimuli but without task engagement), and (3) Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $\Delta t=10$ ms and verify that almost all bins contain no more than one spike. In this dataset, neurons can update in a parallel (non-multipartite) manner.
@@
We calculate EP values across different conditions and different system sizes, $N$. For each combination of size, condition, and recording session, we randomly selected 10 subsets of neurons and estimated ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$. EP is here interpreted as a statistical measure of time-irreversibility, not energetic dissipation.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$) and that the active condition is associated with the largest normalized EP. Fig. 2(b) shows the matrix of inferred parameters $\theta_{ij}^{*}$, grouped by visual brain area. This shows that our method infers the pairwise coupling coefficients that determine temporal irreversibility in a neural system, revealing a clustered organization that is meaningfully aligned with anatomical brain regions.

### v2

@@ Abstract
Figure 1: Disordered nonequilibrium spin model with 1000 spins. (a) Steady-state EP estimates for different inverse temperatures $\beta$. (b) Asymmetry of inferred parameters, shown against the true coupling asymmetries in the model for $\beta=2.5$ ($R^{2}=0.9831$). Estimates are based on $10^{9}$ state transitions sampled by Monte Carlo.
Our observables of interest are time-lagged correlations,
@@
with conjugate parameters $\theta_{ij}$. Because the system has multipartite dynamics and observables, Eq. (3) can be decomposed into $N$ independent problems, each involving only those transitions where spin $i$ changes, improving computational performance for large systems.
Fig. 1(a) shows the actual and inferred EP at different $\beta$ for $N=1000$, $k=6$. EP increases with $\beta$, and all estimators (${\Sigma}_{\bm{g}}^{\textrm{TUR}}$, $\widehat{{\Sigma}}_{\bm{g}}$, ${\Sigma}_{\bm{g}}$) agree in the near-equilibrium regime of small $\beta$. Importantly, ${\Sigma}_{\bm{g}}$ provides a tight bound on EP even in the far-from-equilibrium regime of large $\beta$. The gap between ${\Sigma}_{\bm{g}}$ and $\widehat{{\Sigma}}_{\bm{g}}$ indicates the onset of highly non-Gaussian statistics from $\beta\approx 2$.
As discussed near Eq. (6), ${\Sigma}_{\bm{g}}\approx{\Sigma}$ implies that the trajectory EP $\sigma(\vec{{\bm{x}}})$ can be closely approximated by a linear combination of observables $g_{ij}$. As shown in the SM, our optimal parameters ${\bm{\theta}}^{*}$ can be used to infer the asymmetry of the coupling constants, $\theta_{ij}^{*}-\theta_{ji}^{*}\approx\beta(w_{ij}-w_{ji})$. This relation is verified in Fig. 1(b).
In the SM, we also consider a different family of coupling matrices, for which this model exhibits a nonequilibrium phase transition [6]. We show that our approach accurately infers EP in this regime, a difficult task for existing approximations [53].
@@
Figure 2: (a) Estimated EP per expected number of spikes $R$, in the Neuropixels Visual Behavior dataset for three conditions. Error bars indicate standard error of the mean. (b) A sample of inferred coupling coefficients $\theta^{*}_{ij}$ grouped by visual area. Here we select 200 neurons with highest firing rate from an active trial. To improve visualization, lower triangle shows $\theta_{ji}^{*}\equiv-\theta_{ij}^{*}$ for $i<j$.
Next, as an example application to biological data, we estimate EP in the Visual Behavior Neuropixels dataset [47], containing spike-train recordings from 81 mice across 103 sessions. EP is here interpreted as a statistical measure of time-irreversibility, not energetic dissipation [13]. The dataset includes spiking activity from multiple brain regions, including visual cortical areas (VISp, VISl, VISal, VISrl, VISam and VISpm) and subcortical structures. We analyze data from visual areas during three conditions: active behavior (visual change detection task), passive replay (same stimuli but without task engagement), and Gabor (receptive field characterization with Gabor stimuli and full-field flashes). We discretize spike trains into temporal bins of length $\Delta t=10$ ms and verify that most bins contain no more than one spike. We note that neurons can update in a parallel (non-multipartite) manner.
@@
We calculate EP estimates across different conditions and system sizes $N$. For each combination of size, condition, and recording session, we randomly select 10 subsets of neurons and estimate ${\Sigma}_{\bm{g}}$ for each subset. For improved comparison, we normalize EP in each condition by the expected number of spikes per bin, $R=\sum_{i}\left(1+\langle x_{i,0}\rangle_{\rm st}\right)/2$.
Fig. 2(a) shows that EP grows superlinearly with size (note that $R\propto N$). It also shows that the active condition is associated with the largest normalized EP. Fig. 2(b) illustrates the matrix of inferred parameters $\theta_{ij}^{*}$ for 200 neurons from an active trial, grouped by visual brain area. These parameters can be understood as the coupling coefficients that best describe temporal irreversibility of the in vivo neural dynamics. Notably, they reveal a clustered organization that is aligned with anatomical brain regions.